OpenStudy (anonymous):

The Mecklenberg County Public Schools stipulate that a standard school day will average at least 5 1/2 hr, excluding meal breaks. For the first four days of one school week, bad weather resulted in school days of 4 hr, 6 1/2 hr, 3 1/2 hr, and 6 1/2 hr. How long must the Friday school day be in order to average at least 5 1/2 hr for the week?

OpenStudy (anonymous):

Well this is an algebra question. So let's set it up so "X" is our 5th day. (I am going to use decimals) (4+6.5+3.5+6.5+X)/5 = 5.5 Step 1) Let's multiply by 5 on both sides to get rid of the dividing by 5. 4+6.5+3.5+6.5+X = 27.5 Step 2) Let's add up all of the non x values so we have one value + X. 20.5+X = 27.5 Step 3) Subtract 20.5 from both sides so X is alone. X = 7 There is your answer! The fifth school day must be 7 hours long to achieve an average of 5 1/2hours in one school week.